OpenStudy (anonymous):

y=x4+4x, dx/dt=4 and x+1. how do i find dy/dt?

OpenStudy (tkhunny):

Just write it... \(\dfrac{dy}{dt} = 4x^{3}\cdot \dfrac{dx}{dt} + 4\dfrac{dx}{dt}\) I'm not sure what the "x+1" is. If the problem is not presented correctly, that could make it more difficult.

OpenStudy (anonymous):

oh sorry it was supposed to be x=1

OpenStudy (anonymous):

thank you! i forgot 4(dx/dt)

OpenStudy (anonymous):

its still saying its wrong :/

OpenStudy (perl):



OpenStudy (perl):

is that copied correctly?

OpenStudy (anonymous):

yeah

OpenStudy (perl):